# Neuro Symbolic Knowledge Reasoning for Procedural Video Question Answering with Knowledge Module Learning

### v3

@@ 4 Knowledge Module Learning
Then we train all KMs jointly using the cross-entropy loss over the correct answer index. At test time, we predict the right answer using the argmax of the cosine similarity scores. One of the advantages of KML (auto-program) approach is that, given the relation types, we do not need to manually select any neural modules, or generate programs manually. We use LLMs such as GPT, DeepSeek or Mistral. More implementation details of KML is presented in the supplementary.
5 Experiments
5.1 Experimental Setup
@@
We conduct experiments on the PKR-QA benchmark to evaluate the effectiveness of VLMs and Neurosymbolic methods for procedural knowledge-based question answering. We use NVIDIA A100 GPUs (80GB VRAM) for conducting experiments with VLMs, NVIDIA GeForce RTX 2080 Ti and A5000 GPUs (60GB VRAM) for training KML, and RTX 3090 GPUs (24GB VRAM) for training knowledge graph embedding methods.
VLM Evaluation Settings.
We evaluate VLMs under five different settings to assess their procedural knowledge reasoning ability: (1) Zero-shot: VLMs take a video segment (8 uniformly sampled frames), a question, and options as input; (2) VLM+P.VRL: VLMs take a video segment (8 uniformly sampled frames), a question, options, and the top-5 step/task categories from P.VRL as input; (3) KG-training: VLMs are tuned using the triplet instances of our PKG as questions using LoRA [17] and then evaluate the fine-tuned VLMs with PKR-QA; (4) QA training: VLMs are tuned using LoRA for 100 epochs with 4 randomly sampled frames as input, demonstrating the impact of few-shot fine-tuning. All models are trained using the training set of 1,700 question-answer pairs; (5) KG+QA training: VLMs are tuned using triplet instances following KG-training, and 1,700 question-answer pairs as described in QA training.
Neurosymbolic Methods.
@@
We explore three variants of KML, differing in how the embedding function $x(;\theta_{x})$ is implemented. In KML-F-CLIP, $x(;\theta_{x})$ is a frozen CLIP text encoder. In KML-CLIP, we initialize $x(;\theta_{x})$ with CLIP embeddings and fine-tune its parameters. In KML-Rand, $\theta_{x}$ is learned from scratch (i.e., randomly initialized). We compare our KML against the following NS methods:
@@
KG Embedding Methods: We compare against standard KG embedding models, including TransE [7], TransH [52], and RotatE [46]. After training, we use LLM-generated programs (as in KML) to perform multi-hop reasoning. These models are selected for their support of compositional reasoning. Embedding dimensions are tuned on a validation set, yielding optimal sizes of 256 for TransE, 100 for TransH, and 256 for RotatE. We also include variants with CLIP-initialized entity and relation embeddings to enable direct comparison with CLIP-based KML models.
Modern NS methods: We compare with modern NS methods including ViperGPT [47] that uses the power of LLM and vision models, and MAC [19]. We evaluate MAC on a classification-based VQA task, using a single image frame and a question (without answer choices) as input. The model predicts from 2,079 answer classes aggregated from the dataset. We use GloVe [37] embeddings for text and extract visual features with a pretrained ResNet101 [15]. We tuned the embedding size, MAC hidden size, and the number of MAC layers, selecting the best setup based on validation performance.
Metrics.
Since VLM-generated text may not exactly match the predefined multiple-choice options, we adopt the filtering and MCQ accuracy computation strategy from [63, 27]. We report both overall accuracy and mean accuracy for each model. Accuracy is computed as the average score across all test samples, while mean accuracy (mAcc)s is the average of per-template accuracies, providing equal weight to each traversal template.
5.2 Analysis and Discussion
@@
Zero-shot DeepSeek-VL2 (27.4B) [55] 58.4 55.4
MiniCPM-V (8B) [59] 62.6 59.7
mPLUG-Owl3 (7B) [60] 63.1 60.2
Qwen2.5-VL (7B) [5] 59.6 57.8
VideoChat2-HD (7B) [24] 61.2 58.4
VLM + P.VRL DeepSeek-VL2 (27.4B) [55] 64.5 59.9
MiniCPM-V (8B) [59] 67.4 63.8
mPLUG-Owl3 (7B) [60] 65.5 61.6
Qwen2.5-VL (7B) [5] 69.4 65.8
VideoChat2-HD (7B) [24] 65.5 59.9
KG-training MiniCPM-V (8B) [59] 63.5 61.0
mPLUG-Owl3 (7B) [60] 64.8 61.4
Qwen2.5-VL (7B) [5] 67.3 64.1
QA training MiniCPM-V (8B) [59] 71.1 71.4
mPLUG-Owl3 (7B) [60] 71.8 72.4
Qwen2.5-VL (7B) [5] 73.6 73.4
KG+QA training MiniCPM-V(8B) [59] 72.1 72.1
mPLUG-Owl3 (7B) [60] 73.1 73.8
Qwen2.5-VL (7B) [5] 74.2 73.8
Table 1: Comparison of VLMs in different settings. Underlined scores denote the best-performing method within each setting, while bold scores highlight the best overall.
@@
Table 2: Performance comparison of NS methods.
@@
Table 1 compares the performance of various VLMs across five training and inference settings, revealing several key insights. Providing predicted step and task information from ProceduralVRL (VLM+P.VRL) consistently improves performance over the zero-shot setting. These gains highlight the importance of grounded procedural context in enhancing reasoning, even for strong models like MiniCPM-V and Qwen2.5-VL. Training on KG triplets yields some improvement over zero-shot baselines, though the gains are modest and less consistent, suggesting that aligning symbolic representations with multimodal inputs remains non-trivial. QA-based fine-tuning leads to larger improvements, particularly for MiniCPM-V and Qwen2.5-VL. These gains indicate that VLMs are capable of adapting to task-specific supervision, even when provided with a relatively small number of QA pairs (1,700 samples). The best performance is achieved when combining both KG and QA training, with Qwen2.5-VL reaching 74.2/73.8, indicating that integrating structured procedural knowledge with task-specific examples provides complementary benefits for enhancing procedural understanding in VLMs.
@@
As shown in Table 2, all KML variants outperform all baselines, demonstrating the benefit of executing LLM-generated KG traversal programs with relation-specific KM for the proposed procedural reasoning question answering task. KML achieves the highest performance under KG+QA training using KML-CLIP variant. The performance of KML-Rand is not far from KML-F-CLIP and interestingly, when tuned with KG+QA training the KML-Rand performs better than KML-F-CLIP under the same setting. ViperGPT performs poorly on our benchmark due to the absence of built-in reasoning operators and reliance on predefined program templates that are not suited for the structured reasoning required in procedural knowledge tasks. IGP suffers from combinatorial explosion and poor grounding, often generating irrelevant results from uncertain starting nodes. In contrast, KML handles such uncertainty better by grounding traversal via learned knowledge modules. MAC, originally designed for visual reasoning tasks (e.g., answering questions like “What is to the left of the green box?”), fails in our setting due to its lack of access to external procedural knowledge. Among KG embedding methods, TransE and TransH perform better due to their additive or projective composition, aligning well with our program-based reasoning. RotatE underperforms, likely because its complex-valued, rotation-based composition is less effective for multi-hop reasoning. CLIP-based initialization yields mixed results—improving RotatE but degrading TransE and TransH—indicating varying alignment with visual semantics across models. We also evaluate GPT-4o using top-5 predicted step/task category names from P.VRL. It achieves 71.2% accuracy and 69.0% mean accuracy.
KML’s Interpretability.
One key advantage of KML is its interpretability, as illustrated in the qualitative examples in Table 3 and Table 4. We observe step-by-step reasoning and intermediate interpretations from the learned embeddings, offering insight into the model’s decision process. The output entities represented by the output vectors of each KM seems reasonable and accurate for the given task.
KML’s Generalizability.
To assess KML’s generalizability, we added ten binary relations from the STAR benchmark. KML-F-CLIP achieved 74.9% on Interaction and 76.7% on Feasibility questions, outperforming prior bests of 71.8% [21] and 62.4% [61]. For Sequence and Prediction questions, it scored 57.3% and 49.8%, respectively. We also trained KML-F-CLIP using a GPT-4o-generated KG from 7,687 WikiHow tasks (57,027 steps, 8.7M triplets) capturing tools, actions, objects, and purposes. Evaluated on our PKR-QA  dataset, it achieved 73.9% mean accuracy, rising to 74.9% with KG+QA training, suggesting that while generic KGs help, domain-specific modules remain advantageous.
KML’s Robustness.

### v5

@@ V-B Question Answering suing KML
Then we train all KMs jointly using the cross-entropy loss over answer score distribution where the correct answer is used as the ground truth label for cross-entropy training. At test time, we predict the right answer using the argmax of the cosine similarity scores. One of the advantages of KML (auto-program) approach is that, given the relation types, we do not need to manually select any neural modules, or generate programs manually. We use LLMs such as GPT, DeepSeek or Mistral.
VI Theory
VI-A Theoretical Analysis of KML Learning
In this section, we provide a formal analysis of learning objective of KMs and derive conditions under which the learned mapping satisfies the desired separation property. The ideal outcome of KML learning for relation $R_{k}$ is that the mapped representation $\hat{z}=\phi_{R_{k}}(x)$ separates positives from negatives:
$\min_{y\in Y(x)}s(\hat{z},y)\;>\;\max_{u\in U(x)}s(\hat{z},u).$ (10)
This condition ensures that even the least similar positive tail entity is closer to $\hat{z}$ than the most similar negative entity.
While KML is trained with contrastive loss as in Equation 2. For clarity in this theoretical analysis, we omit the temperature parameter, as it does not affect the ordering arguments.
Let us define
$A:=\sum_{y\in Y(x)}e^{s_{y}},\quad C:=\sum_{u\in U(x)}e^{s_{u}},\quad B:=A+C,$
where $s_{v}:=s(\hat{z},v)$.
Lemma 1 (Loss bound implies relative mass constraint).
If $\mathcal{L}(x)\leq\varepsilon$, then
$C\;\leq\;(e^{\varepsilon}-1)\,A.$
Proof.
From the definition of $\mathcal{L}(x)$,
$\mathcal{L}(x)\leq\varepsilon\;\;\Longleftrightarrow\;\;\frac{A}{A+C}\geq e^{-\varepsilon}.$
Rearranging yields
$C\leq(e^{\varepsilon}-1)\,A.$
Lemma 2 (Bounds on extreme similarities).
The following inequalities hold:
$\min_{y\in Y(x)}s_{y}\;\geq\;\log A-\log|Y(x)|,\qquad\max_{u\in U(x)}s_{u}\;\leq\;\log C.$
Proof.
For the positive set,
$A=\sum_{y\in Y(x)}e^{s_{y}}\;\geq\;|Y(x)|\,e^{\min_{y\in Y(x)}s_{y}},$
which implies
$\min_{y\in Y(x)}s_{y}\geq\log A-\log|Y(x)|.$
For the negative set,
$C=\sum_{u\in U(x)}e^{s_{u}}\;\geq\;e^{\max_{u\in U(x)}s_{u}},$
which implies
$\max_{u\in U(x)}s_{u}\leq\log C.$
Theorem 1 (Sufficient condition for KML separation).
If
$\mathcal{L}(x)\;<\;\log\!\left(1+\frac{1}{|Y(x)|}\right),$ (11)
then the separation condition in Equation 10 holds.
Proof.
From Lemmas 1 and 2,
$\min_{y\in Y(x)}s_{y}-\max_{u\in U(x)}s_{u}\;\geq\;-\log|Y(x)|-\log(e^{\varepsilon}-1),$
where $\varepsilon=\mathcal{L}(x)$. The right-hand side is positive if and only if
$|Y(x)|\,(e^{\varepsilon}-1)<1,$
which is equivalent to
$\varepsilon<\log\!\left(1+\frac{1}{|Y(x)|}\right).$
Under this condition,
$\min_{y\in Y(x)}s_{y}>\max_{u\in U(x)}s_{u},$
and therefore the KML separation condition holds. ∎
Interpretation.
Theorem 1 establishes a concrete link between the value of the KML contrastive loss and the correctness of relational reasoning. When the loss falls below the bound in Equation 11, the learned knowledge module is guaranteed to separate valid tail entities from invalid ones under cosine similarity. This provides an explicit criterion indicating when KML learning is expected to succeed and when separation may fail, offering theoretical insight into the behavior of the model during training.
VI-B Deterministic Composition Error Bound over Traversals
In KML, a reasoning program (or traversal) is a sequence of relation modules. We provide a deterministic error bound on the deviation between the output of the learned Knowledge Modules (KMs) and the output of an ideal (ground-truth) relational operator sequence. The bound explicitly shows how errors accumulate across hops and how Lipschitz continuity controls their propagation. A length-$T$ program is a sequence of relations
$P=(r_{1},r_{2},\ldots,r_{T}).$
For each relation $r_{t}$, KML learns a module $\phi_{r_{t}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ (a small MLP in our implementation). To analyse approximation, we introduce an ideal operator $F_{r_{t}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ that represents the desired mapping induced by the KG relation semantics (e.g., mapping a head-entity representation to a canonical representation of the valid tail-set for that relation).
Given an input state $z_{0}$ (the grounded entity representation), the ideal and learned executions are
$z_{t}$ $=F_{r_{t}}(z_{t-1}),\qquad t=1,\ldots,T,$ (12)
$\hat{z}_{t}$ $=\phi_{r_{t}}(\hat{z}_{t-1}),\qquad t=1,\ldots,T,$ (13)
where $z_{t}$ is the ideal state after $t$ hops and $\hat{z}_{t}$ is the KML-computed state. We define the per-hop (deterministic) deviation using L2 loss as follows:
$\delta_{t}:=\|\hat{z}_{t}-z_{t}\|_{2}.$ (14)
We quantify how well each learned module $\phi_{r_{t}}$ approximates its ideal operator $F_{r_{t}}$ on the unit sphere:
$\varepsilon_{t}:=\sup_{\|u\|_{2}=1}\|\phi_{r_{t}}(u)-F_{r_{t}}(u)\|_{2}.$ (15)
Intuitively, $\varepsilon_{t}$ is the worst-case (deterministic) approximation error of module $r_{t}$ over all unit-norm inputs.
We know each learned module is Lipschitz continuos on the unit sphere by design:
$\|\phi_{r_{t}}(u)-\phi_{r_{t}}(v)\|_{2}\leq L_{t}\|u-v\|_{2},\quad\forall\|u\|_{2}=\|v\|_{2}=1,$ (16)
for some constant $L_{t}\geq 0$ that may depend on the relation $r_{t}$. This is a standard and verifiable property for MLPs: if each layer has bounded spectral norm and the activation is Lipschitz (e.g., $\tanh$ is 1-Lipschitz), then the whole network is Lipschitz with constant bounded by the product of layer norms.
Lemma 3 (One-step error recursion under Lipschitz learned module).).
For each hop $t$,
$\delta_{t}\leq L_{t}\,\delta_{t-1}+\varepsilon_{t}.$ (17)
Proof.
We start from the definition of $\delta_{t}$ and substitute the executions (12)–(13):
$\delta_{t}$ $=\|\hat{z}_{t}-z_{t}\|_{2}$
$=\|\phi_{r_{t}}(\hat{z}_{t-1})-F_{r_{t}}(z_{t-1})\|_{2}.$
We now add and subtract $F_{r_{t}}(\hat{z}_{t-1})$ inside the norm:
$\phi_{r_{t}}(\hat{z}_{t-1})-F_{r_{t}}(z_{t-1})=\underbrace{\big(\phi_{r_{t}}(\hat{z}_{t-1})-F_{r_{t}}(\hat{z}_{t-1})\big)}_{\text{(A) module approximation at }\hat{z}_{t-1}}+$
$\underbrace{\big(F_{r_{t}}(\hat{z}_{t-1})-F_{r_{t}}(z_{t-1})\big)}_{\text{(B) mismatch due to different inputs}}.$
Applying the triangle inequality gives
$\delta_{t}$ $\leq\|\phi_{r_{t}}(\hat{z}_{t-1})-F_{r_{t}}(\hat{z}_{t-1})\|_{2}+\|F_{r_{t}}(\hat{z}_{t-1})-F_{r_{t}}(z_{t-1})\|_{2}.$
At this point, we use the fact that we are analyzing KML under learned-module Lipschitzness. To bound term (B) without requiring Lipschitzness of $F_{r_{t}}$, we express the mismatch through the learned module:
$F_{r_{t}}(\hat{z}_{t-1})-F_{r_{t}}(z_{t-1})\;=$ $\big(F_{r_{t}}(\hat{z}_{t-1})-\phi_{r_{t}}(\hat{z}_{t-1})\big)$ (18)
$+\big(\phi_{r_{t}}(\hat{z}_{t-1})-\phi_{r_{t}}(z_{t-1})\big)$
$+\big(\phi_{r_{t}}(z_{t-1})-F_{r_{t}}(z_{t-1})\big).$
Taking norms and applying triangle inequality again yields
$\|F_{r_{t}}(\hat{z}_{t-1})-F_{r_{t}}(z_{t-1})\|_{2}$ $\leq\|F_{r_{t}}(\hat{z}_{t-1})-\phi_{r_{t}}(\hat{z}_{t-1})\|_{2}$
$\quad+\|\phi_{r_{t}}(\hat{z}_{t-1})-\phi_{r_{t}}(z_{t-1})\|_{2}$
$\quad+\|\phi_{r_{t}}(z_{t-1})-F_{r_{t}}(z_{t-1})\|_{2}.$
Now, by the definition of $\varepsilon_{t}$ in (15), both approximation terms are bounded:
$\|F_{r_{t}}(\hat{z}_{t-1})-\phi_{r_{t}}(\hat{z}_{t-1})\|_{2}\leq\varepsilon_{t},\qquad\|\phi_{r_{t}}(z_{t-1})-F_{r_{t}}(z_{t-1})\|_{2}\leq\varepsilon_{t},$
because $\|\hat{z}_{t-1}\|_{2}=\|z_{t-1}\|_{2}=1$ by assumption. For the middle term, we apply Lipschitzness of the learned module (16):
$\|\phi_{r_{t}}(\hat{z}_{t-1})-\phi_{r_{t}}(z_{t-1})\|_{2}\leq L_{t}\|\hat{z}_{t-1}-z_{t-1}\|_{2}=L_{t}\,\delta_{t-1}.$
Combining these bounds gives
$\|F_{r_{t}}(\hat{z}_{t-1})-F_{r_{t}}(z_{t-1})\|_{2}\leq 2\varepsilon_{t}+L_{t}\,\delta_{t-1}.$
Returning to the earlier inequality for $\delta_{t}$,
$\delta_{t}\leq\underbrace{\|\phi_{r_{t}}(\hat{z}_{t-1})-F_{r_{t}}(\hat{z}_{t-1})\|_{2}}_{\leq\varepsilon_{t}}+\underbrace{\|F_{r_{t}}(\hat{z}_{t-1})-F_{r_{t}}(z_{t-1})\|_{2}}_{\leq 2\varepsilon_{t}+L_{t}\delta_{t-1}},$
we obtain
$\delta_{t}\leq\varepsilon_{t}+(2\varepsilon_{t}+L_{t}\delta_{t-1})=L_{t}\delta_{t-1}+3\varepsilon_{t}.$
Finally, we absorb the constant factor into the definition of $\varepsilon_{t}$ (or define $\tilde{\varepsilon}_{t}:=3\varepsilon_{t}$), which yields the clean recursion (17). $\blacksquare$
Remark (about constants).
The derivation above shows that, under only the Lipschitzness of $\phi_{r_{t}}$, the one-step recursion holds up to a constant factor multiplying $\varepsilon_{t}$; this factor depends on how the “ideal” operator $F_{r_{t}}$ is defined. For clarity in the main theorem, we use the compact form (17) with $\varepsilon_{t}$ interpreted as a uniform bound on the effective one-step approximation error (which may include a small constant factor).
Theorem 2 ((Deterministic composition error bound over a $T$-hop traversal).).
Assume unit-normalized states and (16) for the learned modules. Then the final-state error after executing a length-$T$ program satisfies
$\delta_{T}\leq\left(\prod_{i=1}^{T}L_{i}\right)\delta_{0}+\sum_{t=1}^{T}\left(\prod_{i=t+1}^{T}L_{i}\right)\varepsilon_{t},$ (19)
where by convention $\prod_{i=T+1}^{T}L_{i}:=1$. In particular, if the initial state is exact ($\delta_{0}=0$), then
$\delta_{T}\leq\sum_{t=1}^{T}\left(\prod_{i=t+1}^{T}L_{i}\right)\varepsilon_{t}.$ (20)
Proof.
From Lemma 1 we have the recursion
$\delta_{t}\leq L_{t}\delta_{t-1}+\varepsilon_{t}.$
We expand this inequality iteratively.
Step 1 (expand $\delta_{1}$).
$\delta_{1}\leq L_{1}\delta_{0}+\varepsilon_{1}.$
Step 2 (expand $\delta_{2}$ using the bound on $\delta_{1}$).
$\delta_{2}$ $\leq L_{2}\delta_{1}+\varepsilon_{2}$
$\leq L_{2}(L_{1}\delta_{0}+\varepsilon_{1})+\varepsilon_{2}$
$=(L_{2}L_{1})\delta_{0}+(L_{2})\varepsilon_{1}+\varepsilon_{2}.$
Step 3 (expand $\delta_{3}$).
$\delta_{3}$ $\leq L_{3}\delta_{2}+\varepsilon_{3}$
$\leq L_{3}\big((L_{2}L_{1})\delta_{0}+(L_{2})\varepsilon_{1}+\varepsilon_{2}\big)+\varepsilon_{3}$
$=(L_{3}L_{2}L_{1})\delta_{0}+(L_{3}L_{2})\varepsilon_{1}+(L_{3})\varepsilon_{2}+\varepsilon_{3}.$
Pattern. After $T$ steps, each earlier error $\varepsilon_{t}$ is multiplied by the product of Lipschitz constants of all subsequent modules, because it is propagated through the remaining hops. Thus,
$\delta_{T}\leq\left(\prod_{i=1}^{T}L_{i}\right)\delta_{0}+\sum_{t=1}^{T}\left(\prod_{i=t+1}^{T}L_{i}\right)\varepsilon_{t},$
which is exactly (19). If $\delta_{0}=0$, we obtain (20). $\blacksquare$
Corollary 1 (Simplified bounds for uniform Lipschitzness and uniform approximation error). Assume $L_{t}\leq L$ and $\varepsilon_{t}\leq\varepsilon$ for all $t=1,\ldots,T$, and $\delta_{0}=0$. Then:
  • If $L=1$, the error grows at most linearly with hop length:
    $\delta_{T}\leq T\,\varepsilon.$ (21)
  • If $L\neq 1$, the error is bounded by a geometric series:
    $\delta_{T}\leq\varepsilon\sum_{k=0}^{T-1}L^{k}=\varepsilon\frac{L^{T}-1}{L-1}.$ (22)
  • In the contractive case $L<1$, the error is uniformly bounded as $T\to\infty$:
    $\delta_{T}\leq\frac{\varepsilon}{1-L}.$ (23)
Proof.
Starting from (20) and using $L_{i}\leq L$ and $\varepsilon_{t}\leq\varepsilon$,
$\delta_{T}\leq\sum_{t=1}^{T}\left(\prod_{i=t+1}^{T}L_{i}\right)\varepsilon_{t}\leq\varepsilon\sum_{t=1}^{T}L^{T-t}=\varepsilon\sum_{k=0}^{T-1}L^{k}.$
Evaluating the sum gives (21) when $L=1$ and (22) when $L\neq 1$. If $L<1$, then $\sum_{k=0}^{T-1}L^{k}\leq\sum_{k=0}^{\infty}L^{k}=1/(1-L)$, yielding (23). $\blacksquare$
Interpretation for KML.
Theorem 2 shows that deterministic errors accumulate along a traversal in a structured way: (i) each hop contributes its own approximation error $\varepsilon_{t}$, (ii) errors injected early in the program are amplified by subsequent module Lipschitz constants, and (iii) if the learned modules are near-non-expansive (e.g., $L_{t}\approx 1$ due to bounded activations such as $\tanh$ and regularized weights), then the overall error grows slowly with traversal length. This provides a deterministic stability guarantee for KML execution over multi-hop programs.
VII Experiments
VII-A Experimental resources
@@
We conduct experiments on the PKR-QA benchmark to evaluate the effectiveness of VLMs and Neurosymbolic methods for procedural knowledge-based question answering. We use NVIDIA A100 GPUs (80GB VRAM) for conducting experiments with VLMs, NVIDIA GeForce RTX 2080 Ti and A5000 GPUs (24GB VRAM) for training KML, and RTX 3090 GPUs (24GB VRAM) for training knowledge graph embedding methods.
VII-B KML training details
@@
We explore three variants of KML, differing in how the embedding function $x(;\theta_{x})$ is implemented. In KML-F-CLIP, $x(;\theta_{x})$ is a frozen CLIP text encoder. In KML-CLIP, we initialize $x(;\theta_{x})$ with CLIP embeddings and fine-tune its parameters. In KML-Rand, $\theta_{x}$ is learned from scratch (i.e., randomly initialized). To train KML, we use batch size of 256 triplets with AdamW [loshchilov2017decoupled] optimizer (0.01 learning rate), and 100 training epochs. Programs are generated by GPT-4o by default. For QA training, KML modules are fine-tuned for 100 additional epochs on video-QA pairs (0.001 learning rate). All KML-F-CLIP models employ 2-layer networks (128 hidden dimensions) with CLIP [radford2021learning] (ViT-B/32) text embeddings.
VII-C VLM Evaluation
We evaluate VLMs under five different settings to assess their procedural knowledge reasoning ability: (1) Zero-shot: VLMs take a video segment (8 uniformly sampled frames), a question, and options as input; (2) VLM+P.VRL: VLMs take a video segment (8 uniformly sampled frames), a question, options, and the top-5 step/task categories from P.VRL as input; (3) KG-training: VLMs are tuned using the triplet instances of our PKG as questions using LoRA [hu2021lora] and then evaluate the fine-tuned VLMs with PKR-QA; (4) QA training: VLMs are tuned using LoRA for 100 epochs with 4 randomly sampled frames as input, demonstrating the impact of few-shot fine-tuning. All models are trained using the training set of 1,700 question-answer pairs; (5) KG+QA training: VLMs are tuned using triplet instances following KG-training, and 1,700 question-answer pairs as described in QA training.
VII-D Neurosymbolic Methods
We compare our KML against the following NS methods.
@@ VII-D Neurosymbolic Methods
KG Embedding Methods: We compare against standard KG embedding models, including TransE [bordes2013translating], TransH [wang2014knowledge], and RotatE [sun2019rotate]. After training, we use LLM-generated programs (as in KML) to perform multi-hop reasoning. These models are selected for their support of compositional reasoning. Embedding dimensions are tuned on a validation set, yielding optimal sizes of 256 for TransE, 100 for TransH, and 256 for RotatE. We also include variants with CLIP-initialized entity and relation embeddings to enable direct comparison with CLIP-based KML models.
Modern NS methods: We compare with modern NS methods including ViperGPT [suris2023vipergpt] that uses the power of LLM and vision models, and MAC [hudson2018compositional]. We evaluate MAC on a classification-based VQA task, using a single image frame and a question (without answer choices) as input. The model predicts from 2,079 answer classes aggregated from the dataset. We use GloVe [pennington2014glove] embeddings for text and extract visual features with pretrained ResNet101 [he2016deep]. We tuned the embedding size, MAC hidden size, and the number of MAC layers, selecting the best setup based on validation performance.
VII-E Metrics
Since VLM-generated text may not exactly match the predefined multiple-choice options, we adopt the filtering and MCQ accuracy computation strategy from [yue2023mmmu, lin2023vila]. We report both overall accuracy and mean accuracy for each model. Accuracy is computed as the average score across all test samples, while mean accuracy (mAcc)s is the average of per-template accuracies, providing equal weight to each traversal template.
VII-F Analysis and Discussion
@@
Zero-shot DeepSeek-VL2 (27.4B) [wu2024deepseek] 58.4 55.4
MiniCPM-V (8B) [yao2024minicpm] 62.6 59.7
mPLUG-Owl3 (7B) [ye2024mplug] 63.1 60.2
Qwen2.5-VL (7B) [bai2025qwen2] 59.6 57.8
VideoChat2-HD (7B) [li2023videochat] 61.2 58.4
VLM + P.VRL DeepSeek-VL2 (27.4B) [wu2024deepseek] 64.5 59.9
MiniCPM-V (8B) [yao2024minicpm] 67.4 63.8
mPLUG-Owl3 (7B) [ye2024mplug] 65.5 61.6
Qwen2.5-VL (7B) [bai2025qwen2] 69.4 65.8
VideoChat2-HD (7B) [li2023videochat] 65.5 59.9
KG-training MiniCPM-V (8B) [yao2024minicpm] 63.5 61.0
mPLUG-Owl3 (7B) [ye2024mplug] 64.8 61.4
Qwen2.5-VL (7B) [bai2025qwen2] 67.3 64.1
QA training MiniCPM-V (8B) [yao2024minicpm] 71.1 71.4
mPLUG-Owl3 (7B) [ye2024mplug] 71.8 72.4
Qwen2.5-VL (7B) [bai2025qwen2] 73.6 73.4
KG+QA training MiniCPM-V(8B) [yao2024minicpm] 72.1 72.1
mPLUG-Owl3 (7B) [ye2024mplug] 73.1 73.8
Qwen2.5-VL (7B) [bai2025qwen2] 74.2 73.8
TABLE I: Comparison of VLMs in different settings. Underlined scores denote the best-performing method within each setting, while bold scores highlight the best overall.
@@
TABLE II: Performance comparison of NS methods.
@@
Table I compares the performance of various VLMs across five training and inference settings, revealing several key insights. Providing predicted step and task information from ProceduralVRL (VLM+P.VRL) consistently improves performance over the zero-shot setting. These gains highlight the importance of grounded procedural context in enhancing reasoning, even for strong models like MiniCPM-V and Qwen2.5-VL. Training on KG triplets yields some improvement over zero-shot baselines, though the gains are modest and less consistent, suggesting that aligning symbolic representations with multimodal inputs remains non-trivial. QA-based fine-tuning leads to larger improvements, particularly for MiniCPM-V and Qwen2.5-VL. These gains indicate that VLMs are capable of adapting to task-specific supervision, even when provided with a relatively small number of QA pairs (1,700 samples). The best performance is achieved when combining both KG and QA training, with Qwen2.5-VL reaching 74.2/73.8, indicating that integrating structured procedural knowledge with task-specific examples provides complementary benefits for enhancing procedural understanding in VLMs.
@@
As shown in Table II, all KML variants outperform all baselines, demonstrating the benefit of executing LLM-generated KG traversal programs with relation-specific KM for the proposed procedural reasoning question answering task. KML achieves the highest performance under KG+QA training using KML-CLIP variant. The performance of KML-Rand is not far from KML-F-CLIP and interestingly, when tuned with KG+QA training the KML-Rand performs better than KML-F-CLIP under the same setting. ViperGPT performs poorly on our benchmark due to the absence of built-in reasoning operators and reliance on predefined program templates that are not suited for the structured reasoning required in procedural knowledge tasks. IGP suffers from combinatorial explosion and poor grounding, often generating irrelevant results from uncertain starting nodes. In contrast, KML handles such uncertainty better by grounding traversal via learned knowledge modules. MAC, originally designed for visual reasoning tasks (e.g., answering questions like “What is to the left of the green box?”), fails in our setting due to its lack of access to external procedural knowledge.
Among KG embedding methods, TransE and TransH perform better due to their additive or projective composition, aligning well with our program-based reasoning. RotatE underperforms, likely because its complex-valued, rotation-based composition is less effective for multi-hop reasoning. CLIP-based initialization yields mixed results—improving RotatE but degrading TransE and TransH—indicating varying alignment with visual semantics across models. We also evaluate GPT-4o using top-5 predicted step/task category names from P.VRL. It achieves 71.2% accuracy and 69.0% mean accuracy.
VII-G KML’s Generalizability.
To assess KML’s generalizability, we added ten binary relations from the STAR benchmark [wu2024star]. We use CLIP [radford2021learning] model for entity recognition such as action, relation, object. Specifically, we fine-tune CLIP vision model on the provided annotations of Action Genome dataset and the STAR Benchmark [wu2024star] for visual entity recognition using contrastive learning. We design 10 knowledge modules from the hyper-graphs of the training set of the dataset. KML-F-CLIP achieved 74.9% on Interaction and 76.7% on Feasibility questions, outperforming prior bests of 71.8% [jaiswal2025learning] and 62.4% [yu2023self]. For Sequence and Prediction questions, it scored 57.3% and 49.8%, respectively.
Training with other KGs. We also trained KML-F-CLIP using a GPT-4o-generated KG from 7,687 WikiHow tasks (57,027 steps, 8.7M triplets) capturing tools, actions, objects, and purposes. Evaluated on our PKR-QA  dataset, it achieved 73.9% mean accuracy, rising to 74.9% with KG+QA training, suggesting that while generic KGs help, domain-specific modules remain advantageous.
VII-H KML’s Robustness.